OpenStudy (anonymous):

Please help me people Given any Cartesian coordinates, (x,y), there are polar coordinates (r,θ) with −π/2<θ≤π/2. Find polar coordinates with −π/2<θ≤π/2 for the following Cartesian coordinates: (a) If (x,y)=(3,3) then (r,θ)=( , ) (b) If (x,y)=(13,6) then (r,θ)=( , ) (c) If (x,y)=(−9,−6) then (r,θ)=( , ), (d) If (x,y)=(20,2) then (r,θ)=( , ), (e) If (x,y)=(−4,6) then (r,θ)=( , ), (f) If (x,y)=(0,6) then (r,θ)=( , )

OpenStudy (anonymous):

use the following conversion formulas. \[\Large r=\sqrt{x^2+y^2}\] \[\Large \theta=\tan^{-1} (\frac{y}{x})\] for example for first point x=3 ,y=3 \[\Large r=\sqrt{3^2+3^2}=\sqrt{18}=3\sqrt{2}\] \[\Large \theta=\tan^{-1}(\frac{3}{3})=45^0=\frac{\pi}{4}\]

OpenStudy (anonymous):

i know how to get the first one but the second one is confusing

OpenStudy (anonymous):

never mind i got it